OpenStudy (anonymous):

if h(3) is 7 and g(3) is 1/2 would it be 7/2

OpenStudy (anonymous):

h(3)*g(3)= 7*1/3=7/3

OpenStudy (anonymous):

but it says determine (h/g)(3)

OpenStudy (anonymous):

s means (h/g)(3)=\[\frac{ 7 }{ \frac{ 1 }{ 2 } }=7*2=14\]

OpenStudy (anonymous):

o

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

if it says (f o g)(1) do you multiply

OpenStudy (anonymous):

f((g))

OpenStudy (anonymous):

the question is (h o g)(1) so do i just multiply

OpenStudy (anonymous):

no whats h and g same as the above question ???

OpenStudy (anonymous):

3 and 2

OpenStudy (anonymous):

there wont be any question like that .... it is may be (h*g) is it ur question or texbook question

OpenStudy (anonymous):

a textbook

OpenStudy (anonymous):

(k ○ h)(1) =

OpenStudy (anonymous):

oo one sec ..... @thomaster .... look through pl

OpenStudy (anonymous):

u should have equation given for h(x)and g(x) once refer back to ur text ...

OpenStudy (anonymous):

h is 3 and g is 2

OpenStudy (anonymous):

@thomaster ??

thomaster (thomaster):

I don't understand any of this... what is asked? :s

OpenStudy (anonymous):

(k ○ h)(1)

OpenStudy (anonymous):

k is 3 and h is 2

thomaster (thomaster):

what's the cirlle between K and H for?

OpenStudy (anonymous):

other way around srry

thomaster (thomaster):

circle*

OpenStudy (anonymous):

idk i think it means multiply im not sure

OpenStudy (anonymous):

if its * u r correct that u have to multiply but if its o they should give a equation to substitute ... and get the ans.. ok

thomaster (thomaster):

Well don't know either :P i'm not an expert in mathematics

OpenStudy (anonymous):

anyways did u get it @beezz12

OpenStudy (anonymous):

no

OpenStudy (anonymous):

well i will give u example wait

OpenStudy (anonymous):

1. Given the functions f(x)=5x,g(x)-x^2+1 and , find a.) (f o g) (x) Answer: a.) (f o g)(x)= ( f ( g ) ) ( x ) = ( f ( x ^ 2 +1 ) ) = 5 ( x ^ 2 + 1 ) = 5 x ^ 2 + 1 did u get this .... this is how we do for (f o g)(x)

OpenStudy (anonymous):

@beezz12

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well clear ... now

OpenStudy (anonymous):

so im guessing its the same thing

OpenStudy (anonymous):

no equation so its multiplication ..without equation how can u solve (fog)